OpenStudy (markchernioglo):

The value of a company, in millions of dollars, during its first 10 years increased by 2% each year. The original valuation of the company was 2.1 million dollars. Write a function to represent the value of the company x years after being founded. How much more was the company worth, in millions of dollars, after 6 years than after 2 years? Explain how you solved this problem.

OpenStudy (markchernioglo):

@AloneS one more! :D

OpenStudy (markchernioglo):

@Awolflover1

OpenStudy (danjs):

This is a exponential function, each year you are adding 2% of the total value at the end of that year, so the total adds up and so does the 2% increase over time.

OpenStudy (markchernioglo):

Huh I see... Just not sure how to put it into equation. :/

OpenStudy (danjs):

after one year Value = (Original Value) + (original Value)*0.02 V = 2.1 + 2.1 *0.02 = 2.1 *(1 + .02) V = 2.1 *(1.02) After 2 years, you take another 2%, multiply again V = 2.1 *(1.02)*(1.02) Each year you add another multimple of 0.02, the overall is \[\large V = 2.1*(1.02)^x\]

OpenStudy (markchernioglo):

Ooooh man I see it now! Haha

OpenStudy (danjs):

It is the exponential form y = A*b^x

OpenStudy (markchernioglo):

Yup, thanks Dan, I really appreciat it. :)

OpenStudy (danjs):

welcome